I would like help with the following problem....
Between 1980 and 1990 the population of Columbus grew 12%.
a) Fill in the blank.. If the 1980 population was P, then the 1990 population was _______ my answer is P+(P*.12)
b) The 1990 Population was 632910. Write an equation and find the 1980 population
______________________________ P=______________
my equation P=632910-(P*.12) but now I'm lost

Answer by stanbon(75887)

Between 1980 and 1990 the population of Columbus grew 12%.
a) Fill in the blank.. If the 1980 population was P, then the 1990 population was _______ my answer is P+(P*.12)
1990 Population = 1.12P
---------------------------
b) The 1990 Population was 632910. Write an equation and find the 1980 population
1990 Pop = 1.12 (1980 Pop)
632910 = 1.12 (1980 Pop)
Divide both sides by 1.12 to get:
1980 Pop = 632910/1.12
= 565098.21
You should round this out because you cannot have 0.21 persons.
=565098
Cheers,
Stan H.